OpenStudy (anonymous):

find the slope and the y intercept of the following 4- 1/5y=8x

OpenStudy (anonymous):

1.6x

OpenStudy (anonymous):

8/5 is the slope

OpenStudy (anonymous):

4 - 1/5 y = 8x 4 -8x = 1/5 y Multiply by 5 20 - 40x = y y = -40x + 20 , comapre with y = mx +b m = -40 = slope b=20 = y intercept

OpenStudy (aroub):

-1/5y=8x-4 y=-40x+20 slope=-40 y-intercept=20

OpenStudy (anonymous):



OpenStudy (anonymous):

I know this is the wrong place but this site is soooo coooL!!

OpenStudy (anonymous):



OpenStudy (anonymous):

-8/5 is the slope

OpenStudy (aroub):

no need to do that..

OpenStudy (aroub):

u just multiply the 5 , not divide or idk what u did! :)

OpenStudy (anonymous):

they are tryiung to give u wat y is not the slope